You are building a deck and have enough lumber to make it 88 feet^2. if you want the length of your deck to be 3 feet longer than the width, what should the dimensions of the deck be?
Answer by fractalier(6550)

Call length and width, L and W.
You want A = LW = 88 and L = W + 3...substituting and solving we get
(W + 3)W = 88
W^2 + 3W - 88 = 0
(W + 11)(W - 8) = 0
W = 8
L = 11